2. Suppose that X and Y very inversely in the X equals two when Y equals eight write the function that models the inverse variation
1. Y=16/x
2. Y=10/x
3. Y= 4x
4. Y=6x

Answer by ankor@dixie-net.com(22740)

Suppose that X and Y varies inversely
y = k/x
in the X equals two when Y equals eight
8 = k/2
Find k
k = 8 * 2
k = 16
write the function that models the inverse variation
y = 16/x